OpenStudy (anonymous):

Verify cscAsecA = 2csc2A

OpenStudy (campbell_st):

look at the right hand side 2csc(A) so you have \[2 \times \frac{1}{\sin(2A)}\] and sin(2A) = 2sin(A)cos(A) make the substitution and you get \[2 \times \frac{1}{2ins(A)\cos(A)}...or.... \frac{2}{2\sin(A)\cos(A)}\] hope this helps... there is still a little work to do cancelling the common factor... and then simplifying